Evaluate the double integral by polar coordinates. Integral D 8- 2x² - 2y²dA,


1
Expert's answer
2021-09-28T15:05:40-0400

"\iint(8-2x^2-2y^2)dA\\
=\iint(8-2x^2-2y^2)(r)(d\theta)(dr)\\
=\iint(8-2(x^2+y^2))(r)(d\theta)(dr)"

Put "x=rcos\theta,y=rsin\theta"

"\therefore x^2+y^2=r^2"

So, double integral us reduced to:

"\iint(8-2r^2)(r)(d\theta)(dr)\\
=\iint(8r-2r^3)(d\theta)(dr)\\
=\int(4r^2-\frac{r^4}{2}+c)(d\theta)\\
=4r^3\theta-\frac{r^4\theta }{2}+c\theta+d 
\\"

where "c,d" are constants of integration.